e 2016s tax year 1-18 Tİr Rat e Schedules and Rate Concepts. An examination of the tax rate sched- es for single taxpayers for 2018 (see the inside cover of the text) indicates that the tax is a given dollar amount plus a percentage of taxable income exceeding a particular level. a. Explain how the given dollar amounts are determined . what is his b. Assuming the taxpayer has a taxable income of $50.000 and is single ctax liability for 20182D c. Same facts as (b). What is the taxpayers marginal tax rate? d. Same facts as (b). What is the taxpayers average tax rate? e. Assuming the taxpayer has tax-exempt interest income from municipal bonds of $30,000, what is the taxpayers effective tax rate?
Answer #1
a) The 2018 tax rates for single tax payers are reproduced below:
Single
Taxable Income Tax Rate
$0 – $9,525 10% of taxable income
$9,526 – $38,700 $952.50 plus 12% of the amount over $9,525
$38,701 – $82,500 $4,453.50 plus 22% of the amount over $38,700
$82,501 – $157,500 $14,089.50 plus 24% of the amount over $82,500
$157,501 – $200,000 $32,089.50 plus 32% of the amount over $157,500
$200,001 – $500,000 $45,689.50 plus 35% of the amount over $200,000
$500,001 or more $150,689.50 plus 37% of the amount over $500,000
The given dollar is the amount of tax payable for a person as per the previous
slabs. For instance, in the case given in (b) of the question, the taxable income
is $50000. The person would have to pay tax as per the third slab of $38701 to
$82500.
The slab says that the tax payable is $4453.50 + 22% of the amount over $38700.
Here, the dollar amount of $4453.50 is the total of the tax payable as per the
previous two slabs. It can be worked out as 9525*10%+(38700-9525)*12% = $ 4,453.50
b) Tax liability for 2018 = 4453.50+(50000-38700)*22% = $ 6,939.50
c) Marginal tax rate is that tax rate applicable for the next one dollar, that is for the
50001th dollar. As it also falls in the same tax bracket of 50000, the tax rate would be 22%.
So the marginal tax rate is 22%
d) Average tax rate = Total tax/Taxable income = 6939.50/50000 = 13.88%
e) Effective tax rate = Total/(Taxable income+Exempt) = 6939.50/(50000+30000) = 8.67%